# WWW MIR (2025)

### v1

@@
A Multi-Granularity Multimodal Retrieval Framework for Multimodal Document Tasks
@@
2 Sun Yat-sen University )
@@ 1 Task Description
Our solution, developed by Team LLMHunter, adopts a training-free hybrid framework that integrates multiple off-the-shelf text and vision-language models in a modular design, enabling robust retrieval without task-specific fine-tuning. By combining diverse retrieval signals—such as semantic text similarity, OCR-enhanced image-text alignment, and layout-aware heuristics—our method emphasizes robustness, zero-shot generalization, and ease of deployment. Without relying on task-specific fine-tuning, our approach achieves competitive performance and establishes a strong baseline for multimodal document retrieval.
@@ 3 Methods
Our overall approach follows a multi-stage retrieval framework based on multi-granularity modeling. Given a text or multimodal query, we first encode both queries and document candidates into a unified embedding space across different levels of granularity, including page-level and region-level representations. Retrieval is initially conducted through similarity-based matching, and fusion strategies are employed to integrate multiple retrieval signals. Finally, reranking modules are used to refine the candidate results for enhanced precision.
Both the M2KR and MMDocIR tasks utilize this multi-granularity design philosophy, with M2KR focusing on fine-grained region retrieval and MMDocIR emphasizing full-page retrieval combined with hierarchical reranking. In the following subsections, we present detailed descriptions of our methods for each track.
@@ 3.1 Overall Retrieval Framework
To address the challenges of multimodal document retrieval, we design a comprehensive retrieval framework that unifies multi-granularity representations and multimodal fusion strategies. The framework is structured into two main components: (1) M2KR, focusing on fine-grained region-level retrieval through integrated visual and textual signals; and (2) MMDocIR, which combines full-page retrieval with detailed regional and OCR-based retrieval stages. Both components share a core architecture of multi-stage matching, similarity-based scoring, and reranking modules to ensure accurate and robust retrieval across diverse document types and query formats. The following sections provide detailed descriptions of these two task-specific solutions.
@@ 3.1.1 M2KR: Multimodal Region Retrieval with Fusion
In the M2KR task, we propose a fine-grained multimodal retrieval framework that integrates visual and textual signals at the region level. By leveraging multi-granularity representations, our method effectively captures the detailed layout and semantic information within document pages to enhance retrieval performance.
@@
Input: In the M2KR task, the retrieval process starts by preparing multi-granularity inputs. As is shown in the Figure 1 . For each candidate page, we segment it into multiple regions to obtain region-level representations. Simultaneously, text descriptions for the candidates are generated using the Qwen2.5-VL model, providing complementary semantic information. The user query, which can be in text, image, or multimodal format, is also processed into a unified embedding space.
Process: During retrieval, three types of matching strategies are applied: pure image retrieval based on region embeddings, multimodal retrieval combining query and region features, and pure text retrieval between query and candidate text descriptions. Cosine similarity (CosSim) is calculated for each modality. The resulting relevance scores from the three matching strategies are then fused to jointly consider visual, multimodal, and textual signals.
Output: After score fusion, a reranking module is employed to refine the candidate results. The fusion strategy effectively balances different modalities, while the final rerank step ensures the retrieval results are aligned with the true query intent.
3.1.2 MMDocIR: Full-Page Retrieval with Multistage Ranking
@@
Input: In the MMDocIR task, the retrieval process begins with a user query provided in text form. As is shown in the Figure 2. Candidate document pages are prepared with their full-page images, segmented region images, and OCR-recognized texts to enable multi-granularity retrieval.
Process: First, ColQwen2-7B is used to encode the user query and the full-page image candidates to perform initial full-page retrieval. Simultaneously, we apply GME-7B to conduct two additional retrieval stages: (1) text-to-OCR-text retrieval, where the query is matched against the OCR texts extracted from the pages; and (2) text-to-region-image retrieval, where the query is matched against segmented regional images. Cosine similarity (CosSim) is calculated for each retrieval path. The relevance scores obtained from full-page retrieval, OCR-text retrieval, and regional-image retrieval are then fused to combine global, textual, and regional information.
Output: After score fusion, a Qwen2.5-VL Reranker is employed to perform fine-grained reranking over the fused candidate results. This reranking module further refines retrieval accuracy by leveraging cross-modal understanding, producing the final retrieval output.
3.2 Fusion and Reranking Mechanism
To maximize retrieval accuracy, we employ a combined strategy of multi-source score fusion and reranking. This ensures that both diverse retrieval paths and final semantic alignment are fully leveraged.
@@ 3.2.1 Multi-Source Score Fusion
3.2.2 Reranking with Qwen-VL
After the fusion step, we apply a reranking process using the Qwen2.5-VL model. This reranker performs a binary relevance judgment, predicting whether each candidate is a correct match (labeled as Yes) or not (labeled as No). Candidates classified as Yes are then promoted to the top of the final ranking list. This reranking step ensures that high-confidence matches are prioritized, further refining retrieval accuracy through cross-modal semantic alignment.
@@ 5 Conclusion
In this work, we propose a unified multi-granularity multimodal retrieval framework, and apply it to two document-level tasks: M2KR and MMDocIR. By leveraging hierarchical encoding, modality-aware retrieval strategies, and reranking mechanisms, our approach progressively enhances retrieval accuracy. Experimental results show a clear performance gain from the integration of layout-aware search and reranking modules, with the final system achieving a score of 65.56. This demonstrates the effectiveness and scalability of our framework in complex document retrieval scenarios.

### v2

@@
A Multi-Granularity Retrieval Framework for Visually-Rich Documents
@@
2 School of Intelligent Systems Engineering, Sun Yat-Sen University, Shenzhen, China )
@@ 1 Task Description
Our solution, developed by Team LLMHunter, adopts a training-free hybrid framework that integrates multiple off-the-shelf text and vision-language models in a modular design, enabling robust retrieval without task-specific fine-tuning. By combining diverse retrieval signals—such as semantic text similarity, OCR-enhanced image-text alignment, and layout-aware heuristics—and incorporating a VLM-based candidate filtering module for semantic verification, our method emphasizes robustness, zero-shot generalization, and ease of deployment. Without relying on task-specific training, our approach achieves competitive performance and establishes a strong baseline for multimodal document retrieval.
@@ 3 Methods
Our overall approach follows a multi-stage retrieval framework based on multi-granularity modeling. Given a text or multimodal query, we first encode both queries and document candidates into a unified embedding space across different levels of granularity, including page-level and region-level representations. Retrieval is initially conducted through similarity-based matching, and fusion strategies are employed to integrate multiple retrieval signals. Finally, a vision-language model (VLM)-based candidate filtering module is used to verify and refine the results for enhanced precision.
Both the M2KR and MMDocIR tasks utilize this multi-granularity design philosophy, with M2KR focusing on fine-grained region retrieval and MMDocIR emphasizing full-page retrieval combined with hierarchical validation. In the following subsections, we present detailed descriptions of our methods for each track.
@@ 3.1 Overall Retrieval Framework
To address the challenges of multimodal document retrieval, we design a comprehensive retrieval framework that unifies multi-granularity representations and multimodal fusion strategies. The framework is structured into two main components: (1) M2KR, focusing on fine-grained region-level retrieval through integrated visual and textual signals; and (2) MMDocIR, which combines full-page retrieval with detailed regional and OCR-based retrieval stages. Both components share a core architecture of multi-stage matching, similarity-based scoring, and VLM-based candidate filtering to ensure accurate and robust retrieval across diverse document types and query formats. The following sections provide detailed descriptions of these two task-specific solutions.
@@ 3.1.1 M2KR: Multimodal Region Retrieval with Fusion
In the M2KR task, we propose a fine-grained multimodal retrieval framework that integrates visual and textual signals at the region level. By leveraging multi-granularity representations, our method effectively captures detailed layout and semantic information within document pages to enhance retrieval performance.
@@
Input: The retrieval process starts by preparing multi-granularity inputs. As shown in Figure 1, each candidate page is segmented into multiple regions to obtain region-level representations. Simultaneously, text descriptions for the candidates are generated using the Qwen2.5-VL model, providing complementary semantic information. The user query, which can be in text, image, or multimodal format, is also processed into a unified embedding space.
Process: Three types of matching strategies are applied: pure image retrieval based on region embeddings, multimodal retrieval combining query and region features, and pure text retrieval between query and candidate text descriptions. Cosine similarity (CosSim) is calculated for each modality. The resulting relevance scores from the three matching strategies are then fused to jointly consider visual, multimodal, and textual signals.
Output: After score fusion, a VLM-based verification module is used to assess the semantic alignment between the query and top candidates. This filtering stage promotes high-confidence matches while discarding false positives, ensuring the final results are both precise and semantically faithful to the query.
3.1.2 MMDocIR: Full-Page Retrieval with Multistage Validation
@@
Input: The retrieval process begins with a user query provided in text form. As shown in Figure 2, candidate document pages are prepared with their full-page images, segmented region images, and OCR-recognized texts to enable multi-granularity retrieval.
Process: First, ColQwen2-7B encodes the user query and full-page image candidates to perform initial full-page retrieval. Simultaneously, we apply GME-7B for two additional retrieval paths: (1) text-to-OCR-text retrieval, where the query is matched against the OCR texts extracted from the pages; and (2) text-to-region-image retrieval, where the query is matched against segmented regional images. Cosine similarity (CosSim) is calculated for each retrieval path. The relevance scores obtained from full-page retrieval, OCR-text retrieval, and regional-image retrieval are fused to combine global, textual, and regional information.
Output: After fusion, a VLM-based verification module—powered by the Qwen2.5-VL model—is applied to semantically validate the top-ranked candidates. Rather than merely reordering, this module conducts fine-grained cross-modal verification to confirm the alignment between the query intent and candidate content, ensuring the final results maintain high precision and reliability.
3.2 Fusion and Verification Mechanism
To maximize retrieval accuracy, we employ a combined strategy of multi-source score fusion and vision-language-based verification. This ensures that both diverse retrieval paths and final semantic alignment are fully leveraged.
@@ 3.2.1 Multi-Source Score Fusion
3.2.2 VLM-Based Candidate Verification
After the fusion step, we apply a vision-language-based verification process using the Qwen2.5-VL model. This module performs semantic verification by predicting whether each candidate is a true match (Yes) or not (No). Verified positive candidates are prioritized in the final output, ensuring that only results with strong semantic alignment are retained. This verification step adds an essential layer of cross-modal understanding, refining the retrieval results beyond traditional similarity-based methods.
@@ 5 Conclusion
In this work, we propose a unified multi-granularity multimodal retrieval framework and apply it to two document-level tasks: M2KR and MMDocIR. By leveraging hierarchical encoding, modality-aware retrieval strategies, and vision-language-based candidate verification, our approach progressively enhances retrieval accuracy. Experimental results show a clear performance gain from the integration of layout-aware search and VLM-based verification modules, with the final system achieving a score of 65.56. These results demonstrate the effectiveness and scalability of our framework in complex multimodal document retrieval scenarios.